Paul Sabin organized Sabin Electronics 10 years ago to produce and sell several electronic devices on which he had secured patents. Although the company has been fairly profitable, it is now experiencing a severe cash shortage. For this reason, it is requesting a $500,000 long-term loan from Gulfport State Bank, $100,000 of which will be used to bolster the Cash account and $400,000 of which will be used to modernize equipment. The company's financial statements for the two most recent years follow: Sabin Electronics Comparative Balance Sheet Last Year This Year Assets Current assets Cash Marketable securities Accounts receivable, net Inventory Prepaid expenses $ 70,000 150,000 18,000 300,000 600,000 22,000 480,000 950,000 20,000 Total current assets Plant and equipment, net 1,520,000 1,480,000 1,090,000 1,370,000 Total assets $3,000,000 $2,460,000 Liabilities and Stockholders Equity Liabilities Current liabilities Bonds payable, 12% $ 800,000 $430,000 600,000 600,000 Total liabilities 1,400,000 1,030,000 Stockholders' equity 750,000 850,000 750,000 Common stock, $15 par Retained earnings 680,000 1,430,000 $3,000,000 $2,460,000 Total stockholders' equity 1,600,000 Total liabilities and equity
Sabin ElectroniCS Comparative Income Statement and Reconciliation This Year Last Year Sales Cost of goods sold $5,000,000 $4,350,000 3,450,000 3,875,000 900,000 548,000 Gross margin Selling and administrative expenses 1,125,000 653,000 Net operating income Interest expense 472,000 72,000 352,000 72,000 400,000 120,000 280,000 84,000 Net income before taxes Income taxes (30%) Net income Common dividends 280,000 110,000 196,000 95,000 Net income retained Beginning retained earnings 170,000 680,000 101,000 579,000 Ending retained earnings $ 850,000 $680,000 During the past year, the company introduced several new product lines and raised the selling prices on a number of old product lines in order to improve its profit margin. The company also hired a new sales manager, who has expanded sales into several new territories. Sales terms are 2/10, n/30. All sales are on account. Required I. To assist in approaching the bank about the loan. Paul has asked you to compute the following ratios for both this year and last year: a. The amount of working capital This Year Last Year Working capital
b. The current ratio. (Round your answers to 2 decimal places.) This Year Last Year Current ratio c. The acid-test ratio. (Round your answers to 2 decimal places.) This YearLast Year Acid-test ratio d. The average collection period. (The accounts receivable at the beginning of last year totaled $250,000.) (Round your intermediate calculations and final answers to 1 decimal place. Use 365 days in a year.) This Year Last Year Average collection period days days
e. The average sale period. (The inventory at the beginning of last year totaled $500,000.) (Round your intermediate calculations and final answers to 1 decimal place. Use 365 days in a year.) This Year Last Year Average sale period l:ay: days f. The operating cycle. (Round your answers to 1 decimal place.) This Year Last Year Operating cycle days days g. The total asset turnover. (The total assets at the beginning of last year were $2,420,000) (Round your answers to 2 decimal places.) This Year Last Year Total asset turnover
h. The debt-to-equity ratio. (Round your answers to 3 decimal places.) This YearLast Year Debt-to-equity ratio The times interest earned ratio. (Round your answers to 1 decimal place.) i. his Year Last Year Times interest earned ratio . The equity multiplier. (The total stockholders' equity at the beginning of last year totaled $1,420,000.) Round your answers to 2 decimal places.) This Year Last Year Equity multiplier
a. Present the balance sheet in common-size format. (Round your percentage answers to 1 decimal place (i.e., 0.123 should be entered as 12.3).) Sabin Electronics Common-Size Balance Sheets This Year Last Year Assets Current assets Cash Marketable securitie Accounts receivable, net Inventory Prepaid expenses Total current assets Plant and equipment, net Total assets Liabilities and Stockholders' Equity Liabilities Current liabilities Bonds payable, 12% Total liabilities Stockholders' equity Common stock, $15 par Retained earnings Total stockholders' equity Total liabilities and equity
b. Present the income statement in common-size format down through net income. (Round your percentage answers to 1 decimal place (ie, 0.123 should be entered as 12.3).) Sabin Electronics Common-Size Income Statements This Year Last Year Sales Cost of goods sold Gross margin Selling and administrative expenses Net operating income Interest expense Net income before taxes Income taxes Net income
Answer #1

Please give positive ratings so I can keep answering. Thanks!

Statement of Earnings This Year % Last Year % Ans a- Working Capital
Sales (all on account)      5,000,000.00 100.0%       4,350,000.00 100.0% V Current Assets-Current Liabilities 720,000.00 660,000.00 O=M-N
Cost of goods sold      3,875,000.00 77.5%       3,450,000.00 79.3%
Gross Margin      1,125,000.00 22.50%           900,000.00 20.7% Ans b- Current Ratio
Selling and admin Expenses         653,000.00 13.1%           548,000.00 12.6% Current Assets/Current Liabilities               1.90               2.53 P=M/N
Net operating Income         472,000.00 9.4%           352,000.00 8.1% D
Interest Expense            72,000.00 1.4%             72,000.00 1.7% E Ans c- Acid Test Ratio
Net Income before taxes         400,000.00 8.0%           280,000.00 6.4% Quick Assets/Quick Liabilities               0.69               1.09 R=Q/N
Income tax @ 30%         120,000.00 2.4%             84,000.00 1.9%
Net Income         280,000.00 5.6%           196,000.00 4.5% Ans d- Average collection period (days)
Common Dividends         110,000.00             95,000.00 365/Accounts Receivables Turnover               28.5               23.1 W=365/U
Net Income Retained         170,000.00           101,000.00
Beginning Retained Earnings         680,000.00           579,000.00 Ans e- Average sale period (days)
Ending Retained Earnings         850,000.00           680,000.00 365/Sales Revenue/Average Inventory               56.6               46.1 X=365/(V/T)
Statement of Financial Position Ans f- Total Assets Turnover
Assets Sales Revenue/Average Total Assets               1.83               1.78 Z=V/Y
Current Assets
Cash            70,000.00 2.3%           150,000.00 6.1% Ans g- Debt to Equity Ratio
Marketable securities                          -   0.0%             18,000.00 0.7% Debt/Equity             0.375             0.420 C=A/B
Accounts Receivable (Net)         480,000.00 16.0%           300,000.00 12.2%
Inventory         950,000.00 31.7%           600,000.00 24.4% Ans h- Times Interest Earned
Prepaid Expense            20,000.00 0.7%             22,000.00 0.9% Net operating Income/Interest Expense                 6.6                 4.9 F=D/E
Total Current Assets      1,520,000.00 50.7%       1,090,000.00 44.3% M
Property Plant & Equipment (Net)      1,480,000.00 49.3%       1,370,000.00 55.7% Ans h- Equity Multiplier
Total Assets      3,000,000.00 100.0%       2,460,000.00 100.0% G Total Assets/Total Stockholders' Equity               1.88               1.72 H=G/B
Liabilities & Stockholders' Equity
Liabilities
Current Liabilities         800,000.00 26.7%           430,000.00 17.5% N
Bonds Payable, 12%         600,000.00 20.0%           600,000.00 24.4% A
Total Liabilities      1,400,000.00 46.7%       1,030,000.00 41.9%
Stockholders' Equity
Common Stock. $ 15 par         750,000.00 25.0%           750,000.00 30.5%
Retained Earnings         850,000.00 28.3%           680,000.00 27.6%
Total Stockholders' Equity      1,600,000.00 53.3%       1,430,000.00 58.1% B
Total Liabilities & Stockholders' Equity      3,000,000.00 100%       2,460,000.00 100.0%
Total Current Assets      1,520,000.00       1,090,000.00
Less: Inventory         950,000.00           600,000.00
Less: Prepaid Expense            20,000.00             22,000.00
Quick Assets         550,000.00          468,000.00 Q
This Year Last Year Beginning of Last Year
Accounts Receivable (Net)         480,000.00           300,000.00           250,000.00
Average Accounts Receivable (Net)         390,000.00           275,000.00 S
Inventory         950,000.00           600,000.00           500,000.00
Average Inventory         775,000.00           550,000.00 T
Total Assets      3,000,000.00       2,460,000.00       2,420,000.00
Average Total Assets      2,730,000.00       2,440,000.00 Y
Accounts Receivables Turnover
Sales Revenue/Average Accounts Receivable                   12.82                     15.82 U=V/S